OpenStudy (anonymous):

Help on #'s 14 and 20 please :) thanks

OpenStudy (anonymous):

Thanks

OpenStudy (dumbcow):

#14 is a kite #20: just set the congruent sides equal to each other 2x -3.5 = x + .5 2y-2.8 = y+2

OpenStudy (anonymous):

so X=? and Y=?

OpenStudy (dumbcow):

you tell me, solve each equation by getting the variable by itself on one side

OpenStudy (anonymous):

so, 3 and 3.6?

OpenStudy (dumbcow):

not quite 2x -3.5 = x + .5 subtract x on both sides x -3.5 = .5 add 3.5 on both sides x = 4

OpenStudy (anonymous):

so x=4 and y=4.8?

OpenStudy (dumbcow):

yep :) don't forget to now plug those values back in to find the side lengths